OpenStudy (anonymous):

If x≠-1,x≠1 or x≠3, then what is (x^2-x)/(2x-6)*(x^2-2x-3)/(x^2-1) ?

OpenStudy (ghazi):

\[\frac{ (x^2-x)}{ (2x-6)*(x^2-2x-3) }*(x^2-1)= \frac{ x(x-1) }{ (2x-6)((x-1)^2-2) }*(x-1)(x+1)\]

OpenStudy (anonymous):

so what should i do from here?